JET-COOLED HIGH RESOLUTION INFRARED SPECTRUM OF THE ν_{6} AND 2ν_{2} BANDS OF CH_{2}F_{2}

\underline{THOMAS J. CRONIN}, XIAOLIANG WANG, GREGORY BETHARDY, and DAVID S. PERRY, Department of Chemistry, University of Akron, Akron, OH 44325-3601.

A jet-cooled high resolution infrared spectrum was recorded in a slit-jet apparatus in the asymmetric C-H stretching region of CH_{2}F_{2} between 3002 and 3036 cm^{−1}. Approximately 630 transitions were detected of which about 250 were assigned to the ν_{6} band and about 170 were assigned to the 2ν_{2} overtone at a resolution of ≈ 0.002 cm^{−1}. Seven sub-bands were identified in each vibrational band. The assigned transitions were fitted to a Watson Hamiltonian for an asymmetric rotor (κ = -.932) in the A-reduced I^{r} representation. The fitted upper state constants in cm^{−1}for each band are tabulated below. Numbers in parantheses indicate the precision of the fit to two standard deviations in the last digit.

Parameter  ν_{6}  2ν_{2}
ν_{0}  3014.05028(12) 3026.2297(2)
A′  1.62868(4)  1.63396(6)
B′  0.354165(5)  0.353673(10)
C′  0.308852(3)  0.311833(10)
Δ′_{j} × 10^{−6}  0.417(12)  0.41(9)
Δ′_{jk} × 10^{−5}  -0.16(5)  -0.16(9)
Δ′_{k} × 10^{−4}  0.21(3)  -0.15(4)
δ′_{j} × 10^{−7}  0.93(9)